# On Geometry and Topology of Hessian Manifolds

### v4

@@ 4.2 The Cheng-Yau Metric
Amongst all compact orientable affine manifolds, those being simultaneously radiant and hyperbolic are special, as they admit canonical Hessian metrics. We shall explain this observation in the present section.
@@
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$. Then there exists a unique convex function $u\in C^{\infty}(\Omega)$ satisfying the following properties:
@@
    $\operatorname{det}(\operatorname{Hess}(u))=\operatorname{exp}(2u)^{n}$, and $u(\textup{{x}})\rightarrow\infty$ as $\textup{{x}}\rightarrow\partial\Omega$;
@@
    $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a complete Riemannian metric on $\Omega$;
  3. 3.
    $L^{*}g=g$ for every linear automorphism $L$ of $\Omega$.
@@
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$. The convex function $u\in C^{\infty}(\Omega)$ satisfying the properties listed in Theorem 4.14 is termed the Cheng-Yau potential of $\Omega$.
The differential of a Cheng-Yau potential enjoys one more important property, namely, it is always a unit covector field.
@@
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$, and let $u\in C^{\infty}(\Omega)$ be the Cheng-Yau potential of $\Omega$. Then the holds that $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv 1$.
@@
For any $t>0$, define function $u_{t}\in C^{\infty}(\Omega)$ by $u_{t}(\textup{{x}}):=u(t\cdot\textup{{x}})+\operatorname{log}(t)$. Then straightforward computation yields that $u_{t}$ is also a Cheng-Yau potential of $\Omega$ for every $k>0$. Therefore, by the uniqueness part of Theorem 4.14, we obtain the functional equation
$u(\textup{{x}})-u(t\cdot\textup{{x}})=\operatorname{log}(t)$
@@
$\frac{u(\textup{{x}}+t\cdot\textup{{x}})-u(\textup{{x}})}{t}=-\operatorname{log}(1+t)/t\rightarrow-1$
@@
$\langle\nabla u(\textup{{x}}),\textup{{x}}\rangle\equiv-1.$ (1)
@@
holds for all $\textup{{x}}\in\Omega$. Substitute equation( 2) back into equation( 1), we finally arrive at $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv 1$ as required. ∎
Using Lemma 4.16, we can prove that Cheng-Yau potentials are harmonic functions.
Lemma 4.17.
Let $\Omega$ be a domain in $\mathbb{R}^{d}$, and let $u\in C^{\infty}(\Omega)$. Suppose that $\operatorname{Hess}(u)$ is nowhere degenerated, so that $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a pseudo-Riemannian metric on $\Omega$. Then, the Laplace–Beltrami operator $\Delta^{g}$ of $(\Omega,g)$ satisfies
$\Delta_{g}\phi=\operatorname{tr}(\operatorname{Hess}(u)^{-1}\operatorname{Hess}(\phi))-\frac{1}{2}\langle\nabla\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u))),\operatorname{Hess}(u)^{-1}\nabla\phi\rangle$
for all $\phi\in C^{\infty}(\Omega)$.
Proof.
This is nothing else but a coordinate-dependent and local version of the formula in Lemma 3.33. ∎
Proposition 4.18.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{d}$, and let $u\in C^{\infty}(\Omega)$ be the Cheng-Yau potential of $\Omega$, so that $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a Riemannian metric on $\Omega$. Then $u$ is a harmonic function on $(\Omega,g)$.
Proof.
Since $u$ is a Cheng-Yau potential, we have $2d\cdot u=\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u)))$. Denote by $\Delta^{g}$ the Laplace–Beltrami operator of $(\Omega,g)$. Then by Lemma 4.16 and Lemma 4.17,
$\Delta_{g}u=$ $\operatorname{tr}(\operatorname{Hess}(u)^{-1}\operatorname{Hess}(u))-\frac{1}{2}\langle\nabla\operatorname{log}(\operatorname{det}(\operatorname{Hess}(u))),\operatorname{Hess}(u)^{-1}\nabla u\rangle$
$=$ $\operatorname{tr}(\operatorname{Id})-d\cdot\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle$
$=$ $d-d\cdot 1$
$=$ $0$
as required. ∎
@@
Let $(M,\nabla)$ be an $n$-dimensional compact orientable affine manifold. Suppose that $(M,\nabla)$ is both radiant and hyperbolic. Then, there exists a unique space-like Lagrangian embedding $\eta$ of $(M,\nabla)$ such that $\varkappa=n\eta$, where $\varkappa$ is the first Koszul form of the Hessian manifold $((M,\nabla),\nabla\eta)$.
@@
$n\cdot df=\frac{1}{2}d\operatorname{log}(\operatorname{det}(\operatorname{Hess}(f)))$
descends to $n\cdot\eta=\varkappa$ on $(M,\nabla)$ so that $\eta$ is a space-like Lagrangian embedding of $(M,\nabla)$ and $\varkappa$ is the first Koszul form of $((M,\nabla),\nabla\eta)$. The uniqueness of $\eta$ also follows immediately from the uniqueness of the Cheng-Yau potential. ∎
@@
Let $((M,\nabla),h)$ be a Hessian manifold, and let $\varkappa\in\Omega^{1}(M)$ be its first Koszul form. The metric $h$ is said to be a Cheng-Yau metric if there exists $\eta\in\Omega^{1}(M)$ and a real constant $k>0$ such that $h=\nabla\eta$ and $\varkappa=k\eta$.
4.3 Koszul’s Uncertainty Principle
In this very short section, we introduce a geometric uncertainty principle on compact Hessian manifolds of Koszul type.
Throughout this section, we use Einstein’s summation convention.
We shall first suggest several lemmata, which will turn out to be useful later.
@@ 4.3 Koszul’s Uncertainty Principle
Let $((M,\nabla),h)$ be a Hessian manifold and let $\varkappa\in\Omega^{1}(M)$ be its first Koszul form. Let $A$ be the Amari–Chentsov tensor of $((M,\nabla),h)$. Then, for any vector field $Y$ on $M$, the identity
$\varkappa(Y)=\frac{1}{2}\sum_{i=1}^{n}A(e_{i},e_{i},Y)$
holds in orthonormal frame fields $e_{1},\dots,e_{n}$ of $(M,h)$.
@@
Fix affine coordinates $x^{1},\dots,x^{n}$ of $(M,\nabla)$. Denote $\partial_{i}:=\partial/\partial x^{i}$ for $i=1,\dots,n$. By Jacobi’s formula for the derivative of a determinant, we have that
$2\varkappa_{k}=\partial_{k}(\operatorname{log}(|h|))=\operatorname{tr}(h^{-1}\partial_{k}h)=h^{ij}\partial_{k}h_{ij}$
where $|h|$ is the determinant of the matrix of which $(i,j)$-entry is $h_{ij}$. Moreover, by definition, we have $A_{ijk}=\partial_{k}h_{ij}$ for all $i,j,k\in\{1,\dots,n\}$. By computing the trace of $Y^{k}A_{ijk}$, we obtain
$2\varkappa_{k}Y^{k}=h^{ij}\partial_{k}h_{ij}Y^{k}=h^{ij}A_{ijk}Y^{k}$
as required. ∎

### v6

@@ 3.3 The Cheng-Yau Metric
Amongst all compact affine manifolds, those being hyperbolic are special, as they admit canonical Hessian metrics. We shall explain this observation in the present section.
@@
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$. Then, for any positive real number $\Lambda\in\mathbb{R}$, there exists a unique convex function $u\in C^{\infty}(\Omega)$ satisfying the following properties:
@@
    $\operatorname{det}(\operatorname{Hess}(u))=\operatorname{exp}(2\Lambda u)$;
@@
    $u(\textup{{x}})\rightarrow\infty$ as $\textup{{x}}\rightarrow\partial\Omega$;
so that in particular $g=(\partial_{i}\partial_{j}u)dx^{i}\otimes dx^{j}$ is a complete Riemannian metric on $\Omega$.
@@
For a regular convex cone $\Omega$ in $\mathbb{R}^{n}$ and a positive real number $\Lambda\in\mathbb{R}$, the convex function $u\in C^{\infty}(\Omega)$ satisfying the properties listed in Theorem 3.24 is termed the Cheng-Yau potential of $\Omega$ with cosmological constant $\Lambda>0$.
Example 3.26.
Denote by $Q(x_{1},\dots,x_{n};y):=y^{2}-(x_{1}^{2}+\cdots+x_{n}^{2})$ the Lorentzian quadratic form on $\mathbb{R}^{1,n}$. The Cheng-Yau potential for the right circular cone $\Omega:=\{(x_{1},\dots,x_{n};y)\in\mathbb{R}^{n}\times\mathbb{R}:Q(x_{1},\dots,x_{n};y)>0\}$ is
$u(x_{1},\dots,x_{n};y)=-\frac{n}{2\Lambda}\operatorname{log}(\Lambda\cdot Q(x_{1},\dots,x_{n};y)/n)$
where $\Lambda>0$ is the cosmological constant.
Proposition 3.27.
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$, and let $u\in C^{\infty}(\Omega)$ be a Cheng-Yau potential of $\Omega$. Then $u$ satisfies $du=L^{*}du$ for every linear automorphism $L$ of $\Omega$.
Proof.
See [2], or see [6] for the original proof. ∎
The differential of a Cheng-Yau potential enjoys one more important property, namely, it is always a covector field of constant length.
@@
Let $\Omega$ be a regular convex cone in $\mathbb{R}^{n}$, and let $u\in C^{\infty}(\Omega)$ be the Cheng-Yau potential of $\Omega$ with cosmological constant $\Lambda>0$. Then $u\in C^{\infty}(\Omega)$ satisfies $\langle\nabla u(\textup{{x}}),\textup{{x}}\rangle\equiv-n/\Lambda$, and
$\operatorname{Hess}(u)^{-1}\nabla u(\textup{{x}})=-\textup{{x}}$
for all $\textup{{x}}\in\Omega$, that is, $\operatorname{Hess}(u)^{-1}\nabla u$ is the standard Euler vector field of $\Omega$. In particular, it holds that $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv n/\Lambda$.
@@
For any $t>0$, define function $u_{t}\in C^{\infty}(\Omega)$ by $u_{t}(\textup{{x}}):=u(t\cdot\textup{{x}})+n\cdot\operatorname{log}(t)/\Lambda$. Then straightforward computation yields that $u_{t}$ is also a Cheng-Yau potential of $\Omega$ for every $t>0$. Therefore, by the uniqueness part of Theorem 3.24, we obtain the functional equation
$u(\textup{{x}})-u(t\cdot\textup{{x}})=n\cdot\operatorname{log}(t)/\Lambda$
@@
$\frac{u(\textup{{x}}+t\cdot\textup{{x}})-u(\textup{{x}})}{(n/\Lambda)t}=-\operatorname{log}(1+t)/t\rightarrow-1$
@@
$\langle\nabla u(\textup{{x}}),\textup{{x}}\rangle\equiv-n/\Lambda.$ (1)
@@
holds for all $\textup{{x}}\in\Omega$. Substitute equation( 2) back into equation( 1), we finally arrive at $\langle\nabla u,\operatorname{Hess}(u)^{-1}\nabla u\rangle\equiv n/\Lambda$ as required. ∎
@@
Let $(M,\nabla)$ be a compact hyperbolic affine manifold. Then, for any positive real number $\Lambda\in\mathbb{R}$, there exists a unique space-like Lagrangian embedding $\eta\in\Omega^{1}(M)$ of $(M,\nabla)$ such that $\varkappa=\Lambda\eta$, where $\varkappa$ is the first Koszul form of the Hessian manifold $((M,\nabla),h)$ and here $h=\nabla\eta$.
@@
$\Lambda\cdot df=\frac{1}{2}d\operatorname{log}(\operatorname{det}(\operatorname{Hess}(f)))$
descends to $\Lambda\cdot\eta=\varkappa$ on $(M,\nabla)$ so that $\eta$ is a space-like Lagrangian embedding of $(M,\nabla)$ and $\varkappa$ is the first Koszul form of $((M,\nabla),\nabla\eta)$. The uniqueness of $\eta$ also follows immediately from the uniqueness of the Cheng-Yau potential. ∎
@@
Let $((M,\nabla),h)$ be a Hessian manifold, and let $\varkappa\in\Omega^{1}(M)$ be its first Koszul form. The metric $h$ is said to be a Cheng-Yau metric with cosmological constant $\Lambda>0$ if $\nabla\varkappa=\Lambda\cdot h$. A Hessian manifold $((M,\nabla),h)$ is termed a Cheng-Yau manifold if $h$ is a Cheng-Yau metric.
We observe a property of Cheng-Yau metrics that their $1$-form potentials are self-dual under Legendre transformation.
Proposition 3.31.
Let $((M,\nabla),h)$ be a Cheng-Yau manifold with cosmological constant $\Lambda>0$. Let $\eta$ be the $1$-form potential of $((M,\nabla),h)$, and $\eta^{*}$ its Legendre transform. Then $\eta^{\sharp}$ is the unique Euler field of $(M,\nabla)$, and it holds that $\eta^{*}=-\eta$, where $\sharp$ is the musical isomorphism.
Proof.
For the first assertion, it suffices to lift $\eta\in\Omega^{1}(M)$ to the universal covering and apply Lemma 3.28. The uniqueness follows from Corollary 3.9.
Again by Lemma 3.28, we have
$\eta(\eta^{\sharp})=|\eta|^{2}=\operatorname{dim}(M)/\Lambda$
is a constant function. By Cartan’s magic formula and the definition of Legendre transformation, we have
$\eta^{*}=d(\eta(-\eta^{\sharp}))-\eta=-\eta$
as required. ∎
4 Fibration and Splitting Theorems
4.1 A Naive Attempt via Tischler’s Construction
Before we delve deeply into the structure theorems of Hessian manifolds, we first suggest here several naive observations on splitting of Hessian manifolds, which all relies on the following fundamental result of Tischler on fibration over circles.
Theorem 4.1 (Tischler, 1970[25]).
Proposition 4.2.
Proof.
Let $\varkappa$ be the first Koszul form of $((M,\nabla),h)$, which is closed by definition. If $\varkappa=0$, then $(M,h)$ is a flat Riemannian manifold by Theorem 2.40. We therefore assume $\varkappa\neq 0$. But then since $\bar{\nabla}\varkappa=0$ by Proposition 2.39, we have that $\varkappa$ is nowhere vanishing by parallel transport.
The desired result now follows from Theorem 4.1. ∎
We now can confirm Chern’s conjecture in our special case:
Corollary 4.3.
The Euler characteristic of a compact Hessian manifold is zero.
Proof.
By Proposition 4.2, $M$ is homeomorphic to a mapping torus or a Bieberbach manifold, and hence $\chi(M)=0$ by the multiplicativity of Euler characteristic. ∎
By showing that the $1$-form potential is also nowhere vanishing, we give an alternative proof of the fibration theorem.
Proposition 4.4.
The $1$-form potential of a compact Hessian manifold of Koszul type is a nowhere vanishing closed 1-form.
Proof.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type, and let $\eta\in\Omega^{1}(M)$ be its $1$-form potential.
View $\eta\colon M\rightarrow T^{*}M$ as an embedding. Then, $\eta$ vanishes at $p\in M$ if and only if $p$ is an intersection point of $\eta\colon M\rightarrow T^{*}M$ and the zero section $\sigma$ of the cotangent bundle of $M$. Since $\nabla\eta=h>0$, each intersection point of $\eta$ and $\sigma$ is a nondegenerate transverse intersection, and the intersection numbers at those intersection points are all positive. But, on the other hand, the sum of the intersection numbers at the intersection points of $\eta$ and $\sigma$ is precisely
$\langle e(T^{*}M),[M]\rangle=\chi(M)=0$
by Proposition 4.3. This proves that the sections $\eta$ and $\sigma$ must have no intersection, i.e. $\eta$ is a nowhere vanishing closed 1-form on $M$. ∎
Corollary 4.5.
Let $((M,\nabla),h)$ be a compact Hessian manifold of Koszul type. Then there exists a fiber bundle $\varphi\colon M\rightarrow\mathbb{S}^{1}$ with connected fiber. Moreover, if $M$ is oriented, then the fiber of $\varphi\colon M\rightarrow\mathbb{S}^{1}$ is also oriented.
Proof.
The desired result follows immediately from Theorem 4.1 and Proposition 4.4. ∎
4.2 Rational Approximation of Affine Structures
Making use of the existence of Cheng-Yau metric, we now introduce an important property of compact hyperbolic affine manifolds.
Definition 4.6.
We recall that, by Proposition 2.36, when the affine manifold under consideration is equipped with a Hessian metric, its first Koszul form represents precisely its canonical class.
Compact orientable Hessian manifolds with rational canonical classes are particularly well-behaved, as their mapping tori structures are compatible with their Hessian metrics.
Theorem 4.7.
Let $((M,\nabla),h)$ be a compact oriented Hessian manifold with non-zero rational canonical class. Then there exists a Riemannian submersion $f\colon M\rightarrow\mathbb{S}^{1}$ with connected and oriented fiber $F$, such that the monodromy of $f\colon M\rightarrow\mathbb{S}^{1}$ is an isometry of $(F,h|_{F})$ and in particular is of finite order.
Proof.
Let $\varkappa$ be the first Koszul form and $\bar{\nabla}$ the Levi-Civita connection of $((M,\nabla),h)$. Since $\bar{\nabla}\varkappa=0$ by Proposition 2.39 and $\varkappa\neq 0$ by assumption, we have that $\varkappa$ is nowhere vanishing via parallel transport.
Since $\varkappa]\in\mathrm{H}^{1}(M;\mathbb{Q})$, the rational number
$m:=\operatorname{min}\{k\in\mathbb{R}:k>0,k[\varkappa]\in\mathrm{H}^{1}(M;\mathbb{Z})\}$
is well-defined. Fix a base point $q$ on $M$, and define the smooth mapping $f\colon M\rightarrow\mathbb{S}^{1}$ via
$f(p):=\operatorname{exp}\left(2\pi\sqrt{-1}\int_{\gamma}m\varkappa\right)$
where $\gamma\colon[0,1]\rightarrow M$ is a smooth curve with $\gamma(0)=q$ and $\gamma(1)=p$. Since the periods of $m[\varkappa]$ are integers, the map $f$ is well-defined. Moreover, since $\varkappa$ is nowhere vanishing, we have that $f$ is a submersion. By Ehresmann’s fibration theorem, we have that $f\colon M\rightarrow\mathbb{S}^{1}$ is a fiber bundle. Denote by $F$ the fiber of $f\colon M\rightarrow\mathbb{S}^{1}$.
Identifying each homotopy class in $\pi_{1}(\mathbb{S}^{1})$ with its winding number, it is readily seen that the induced homomorphism $f_{*}\colon\pi_{1}(M)\rightarrow\pi_{1}(\mathbb{S}^{1})=\mathbb{Z}$ satisfies
$f_{*}([\gamma])=\int_{\gamma}m\varkappa$
for all $\gamma]\in\pi_{1}(M)$. Since $\varkappa\neq 0$, we have that $\operatorname{Im}(f_{*})\neq 0$. Therefore, there exists a positive integer $d\in\mathbb{Z}$ such that $\operatorname{Im}(f_{*})=d\mathbb{Z}$. But by the minimality of the factor $m$, we must have that $d=1$ and hence $\operatorname{Im}(f_{*})=\mathbb{Z}$. Now, since $M$ is connected and $f_{*}$ is surjective, the long exact sequence
Moreover, since $f\colon M\rightarrow\mathbb{S}^{1}$ is a submersion between compact differentiable manifolds, we have that $F$ is also a compact differentiable manifold. Also, the fiber $F$ is oriented as $M$ is.
Denote by $\varphi$ the monodromy of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$. Let $\pi\colon\tilde{M}\rightarrow M$ be the universal covering of $M$. Then there exists a complete Riemannian metric $g$ on $\mathbb{R}^{n}$ such that $(\tilde{M},\pi^{*}h)$ is isometric to the product of Riemannian manifolds $(\mathbb{R}^{n},g)$ and $(\mathbb{R},dt^{2})$. Denote by $\hat{\varphi}\colon\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ the lifting of $\varphi\colon F\rightarrow F$ to the universal covering spaces. Since $\pi_{1}(M)=\pi_{1}(F)\rtimes_{\varphi}\mathbb{Z}$ acts on $(\mathbb{R}^{n}\times\mathbb{R},g\oplus dt^{2})$ by isometry, in particular the smooth mapping
is an isometry. Therefore we conclude that $\hat{\varphi}$ and hence $\varphi$ are isometries. Furthermore, since $f\colon M\rightarrow\mathbb{S}^{1}$ lifts to the coordinate projection $\mathbb{R}^{n}\times\mathbb{R}\rightarrow\mathbb{R}$ on universal covering spaces, we conclude that $f$ is a Riemannian submersion.
Denote by $i\colon F\rightarrow M$ the inclusion map of the fiber of the fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$. Consider the subgroup $A$ of the isometry group $G$ of $(F,i^{*}h)$ generated by $\varphi$, and denote by $\bar{A}$ its closure in $G$. Then it is readily seen that $\bar{A}$ is a compact abelian subgroup of $G$, and hence the identity component of $\bar{A}$ is a torus, Therefore, $\varphi$ is isotopic in $G$ to a group element of finite order. ∎
We first introduce the moduli space of radiant affine structures on a differentiable manifold.
@@ 4.2 Rational Approximation of Affine Structures
Let $M$ be an $n$-dimensional differentiable manifold and $G:=\pi_{1}(M)$ its fundamental group. Then the moduli space of radiant affine structures on $M$ is a real analytic subvariety of the character variety $\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))/\operatorname{GL}_{n}(\mathbb{R})$.
@@
The moduli space of flat connections on the frame bundle of $M$ is the character variety $\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))/\operatorname{GL}_{n}(\mathbb{R})$, whilst the torsion-free condition imposes a system of transcendental equations on $\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))/\operatorname{GL}_{n}(\mathbb{R})$. ∎
To proceed, we adapt a theorem of J.-L.Koszul to our setting.
Theorem 4.9 (Koszul, 1968[19]).
Let $(M,\nabla)$ be a compact hyperbolic affine manifold. Then the collection of hyperbolic affine structures on $M$ is an non-empty open subset of the moduli space of radiant affine structures on $M$.
Applying Theorem 4.9 together with group cohomologies with Lie algebra coefficients, we can prove the following density or approximation theorem:
Definition 4.10.
For an affine manifold $(M,\nabla)$, a sequence $\{\nabla^{i}\}_{i=1}^{\infty}$ of flat torsion-free connections on the tangent bundle $TM\rightarrow M$ is said to be a convergent sequence of differential operators with limit $\nabla$, if for any vector fields $X,Y$ on $M$, the sequence $\{\nabla^{i}_{X}Y\}_{i=1}^{\infty}$ of vector fields converges uniformly to $\nabla_{X}Y$.
Theorem 4.11.
Let $(M,\nabla)$ be a compact hyperbolic affine manifold. Then there exists a sequence $\{\nabla^{i}\}_{i=1}^{\infty}$ of hyperbolic affine structures converging to $\nabla$, such that for every integer $i\geq 1$ the canonical class of $(M,\nabla^{i})$ is rational.
Proof.
For simplicity, denote by $G:=\pi_{1}(M)$ the fundamental group of $M$. Let $\mathfrak{X}$ be the moduli space of radiant affine structures on $M$. By Lemma 4.8, $\mathfrak{X}$ is a real analytic subvariety of the character variety $\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))/\operatorname{GL}_{n}(\mathbb{R})$. By algebraic geometry, the singular locus of $\mathfrak{X}$ is of positive codimension. Also, recall that by Theorem 4.9 the collection of hyperbolic affine structures on $M$ is an non-empty open subset of $\mathfrak{X}$. Therefore, without loss of generality, we may assume that the gauge equivalent class $A:=[\nabla]$ of $\nabla$ is a smooth point of $\mathfrak{X}$.
The $\operatorname{GL}_{n}(\mathbb{R})$-invariant epimorphism
$\varphi\colon\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))\rightarrow\operatorname{Hom}(G,\mathbb{R}^{\times}),\quad\rho\mapsto(g\mapsto\operatorname{det}(\rho(g)))$
induces a surjective regular morphism
$\hat{\varphi}\colon\operatorname{Hom}(G,\operatorname{GL}_{n}(\mathbb{R}))/\operatorname{GL}_{n}(\mathbb{R})\rightarrow\operatorname{Hom}(G,\mathbb{R}^{\times}),\quad[\rho]\mapsto\varphi(\rho).$
Since the induced homomorphism
of the trace map
is surjective as the short exact sequence
[…]
Let $h$ be the Cheng-Yau metric of $(M,\nabla)$ with cosmological constant $\Lambda=1$, and let $\eta$ be its $1$-form potential. Recall from Proposition 3.31 that the Legendre transform $\eta^{*}$ of $\eta$ is precisely $\eta$, and hence $\eta^{\sharp}$ is the Euler vector field of $(M,\nabla)$ where $\sharp$ is the musical isomorphism induced by $h$. Consider the $(1,2)$-tensor $B$ defined by
$B(X,Y):=\eta(X)Y+\eta(Y)X-h(X,Y)\eta^{\sharp}$
for all vector fields $X,Y$ on $M$. Since $B(X,Y)=B(Y,X)$ and $(\nabla_{X}B)(Y,Z)=(\nabla_{Y}B)(X,Z)$ for all vector fields $X,Y,Z$ on $M$ by the very construction of $B$, it is readily seen that $B\in T_{A}(\mathfrak{X})$ is an infinitesimal deformation of hyperbolic affine structures on $M$. Moreover, a computation yields that
$\operatorname{tr}_{h}(B)=\eta^{*}+(n+1)\eta=\operatorname{dim}(M)\eta$
is closed but not exact by Proposition 2.8.
[…]
By restricting $\hat{\varphi}$ to a sufficiently small neighborhood of $A$ in $\mathfrak{X}$, the desired result follows immediately from the open mapping theorem for submersions and the density of rational cohomology classes in de Rham cohomology groups. ∎
Theorem 4.11 has profound consequence, which significantly strengthen the previous Corollary 4.5.
Corollary 4.12.
Let $(M,\nabla)$ be a compact oriented hyperbolic affine manifold. Then there exists a fiber bundle $f\colon M\rightarrow\mathbb{S}^{1}$ with connected and oriented fiber, such that the monodromy of the mapping torus $f\colon M\rightarrow\mathbb{S}^{1}$ is periodic.
Proof.
By Corollary 3.22 and Theorem 2.40, the canonical class $K$ of $(M,\nabla)$ is non-trivial. By Theorem 4.11, we may assume $K$ is a non-trivial integral cohomology class. Now it suffices to apply Theorem 4.7 to $((M,\nabla),h)$ for the Cheng-Yau metric $h$ of $(M,\nabla)$. ∎
4.3 Convex Decomposition Lemma
We shall now prove yet another splitting theorem for compact Hessian manifolds, showing that up to a finite covering every compact Hessian manifold is a product of a hyperbolic affine manifold with a flat torus.
We start by introducing an easy observation: